# Accurate BGV Parameters Selection: Accounting for Secret and Public Key Dependencies in Average-Case Analysis

### v2

@@ 6 Conclusions
In this work, we focused on improving the current average-case approach for tracking the error and its growth under homomorphic operations. In particular, we demonstrated that the primary reason for the underestimation of noise in existing average-case methods lies in the fact that the coefficients of the error polynomial are treated as independent. However, to obtain accurate bounds that never underestimate the actual values observed in practice, it is crucial to account for the dependencies among these coefficients.
Accurate noise estimation plays a key role in the proper selection of scheme parameters, especially the ciphertext modulus $q$, to ensure correctness, security, and improved efficiency. To evaluate the effectiveness of our analysis, we applied our results to specific homomorphic circuits, comparing our estimations with both experimental data and those obtained from the current average-case approach. The circuits were implemented using the MAGMA language.
Our findings show that the proposed method effectively overcomes the limitations of existing average-case techniques, particularly their tendency to underestimate noise variance. Furthermore, in the context of initial parameter selection, the high accuracy of our estimations, consistently close to experimental results, represents a significant improvement over the worst-case methods that currently dominate state-of-the-art parameter generation. This leads to tighter bounds and thus enables more efficient configurations of the scheme, which is especially critical for promoting the practical adoption of homomorphic encryption, and in particular of the BGV scheme, in real-world applications.
Future work will include extending our approach to other cryptographic schemes, such as CKKS, as well as studying scenarios in which ciphertexts are generated in a dependent manner. We are currently developing a comprehensive framework for modulus selection that is fully aligned with the new approach introduced in this paper.
@@ Acknowledgment
The third author was partially supported by the Italian Ministry of University and Research in the framework of the Call for Proposals for scrolling of final rankings of the PRIN 2022 call - Protocol no. 2022RFAZCJ and by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
@@ Appendix 0.A Proof of Lemma 1
$b^{\mathsf{mul}}_{\mu}({\iota})=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}b_{\mu_{1}}({\iota_{1}})b^{\prime}_{\mu_{2}}({\iota_{2}}).$
@@
$\mathbb{E}[b^{\mathsf{mul}}_{\mu}({\iota})]$ $=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})b^{\prime}_{\mu_{2}}({\iota_{2}})]$
$=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})]\mathbb{E}[b^{\prime}_{\mu_{2}}({\iota_{2}})]=0,$
@@
$\nu^{\prime}=\frac{q_{l^{\prime}}}{q_{l}}(\nu+\nu_{\mathsf{ms}})\quad\text{ where}\quad\nu_{\mathsf{ms}}=\delta_{0}+\delta_{1}s$
@@ Appendix 0.B Proof of Lemma 2
Recall that, given two random variables X and Y, the following properties hold:
[…]
@@
[…]
where the second property can be generalized for $k$ random variables $\{X_{i}\}_{i=0}^{k}$ as follows
@@
where the covariances vanishes based on property 1.
Thus, the following equality holds
@@
This results in
$\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})=\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})(\mathsf{Var}(s^{\iota}|_{i-j})+\mathbb{E}[s^{\iota}|_{i-j}]^{2}).$
In addition, for a random variable $X$, it holds that
$\mathsf{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}.$
Therefore,

### v3

@@ 8 Conclusions
In this work, we introduce a new approach to average cases that can accurately estimate the error arising from homomorphic operations, especially multiplication. The main novelty of our method lies in the definition of error bounds that account for the dependencies between the critical quantities of multiplied ciphertexts, arising from both the public and private keys. We believe that this is precisely what enables our approach to overcome the typical underestimations observed in current average-case analyses, thereby providing accurate bounds on the error variance without ever underestimating it.
Furthermore, this work aims to provide general guidelines for studying noise growth in circuits independently of the specific implementation of the homomorphic encryption library employed. Based on these estimates, we present a method to select ciphertext moduli appropriately in a generic circuit, demonstrating that the new solution facilitates a significant reduction in their size. This leads to improved efficiency compared to current techniques and compared to the moduli employed in major homomorphic encryption libraries, with no loss of security and proper scheme functioning.
Finally, we provide theoretical and empirical evidence supporting the validity of average-case approaches in the study of homomorphic schemes such as BGV. Precisely, we show that a proper choice of moduli together with application of modulus switching on a systematic basis in BGV allow error distributions to be well approximated by Gaussian distributions. This confirms the validity of such approaches and highlights their practical relevance in improving efficiency, thereby contributing to the potential widespread adoption of these schemes in real-world applications.
@@ Acknowledgment
The third and fourth authors were partially supported by the Italian Ministry of University and Research in the framework of the Call for Proposals for scrolling of final rankings of the PRIN 2022 call - Protocol no. 2022RFAZCJ. The third author acknowledges project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
@@ Appendix 0.A Proof of Lemma 1
$b^{\mathsf{mul}}_{\mu}({\iota})|_{i}=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\sum_{j}b_{\mu_{1}}({\iota_{1}})|_{j}\,b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i-j}.$
@@
$\mathbb{E}[b^{\mathsf{mul}}_{\mu}({\iota})|_{i}]$ $=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\sum_{j}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})|_{j}\,b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i-j}]$
$=\sum_{{\iota_{1}}+{\iota_{2}}=\iota}\sum_{\mu_{1}+\mu_{2}=\mu}\sum_{j}\mathbb{E}[b_{\mu_{1}}({\iota_{1}})|_{j}]\mathbb{E}[b^{\prime}_{\mu_{2}}({\iota_{2}})|_{i-j}]=0,$
@@
$\nu^{\prime}=\frac{q_{l^{\prime}}}{q_{l}}(\nu+\nu_{\mathsf{ms}})\quad\text{ where}\quad\nu_{\mathsf{ms}}=\delta_{0}+\delta_{1}s\,.$
@@ Appendix 0.B Proof of Lemma 2
Thus, given two random variables $X$ and $Y$, the following properties hold:
$\text{a. \hskip 5.69046pt}\mathsf{Var}(XY)=(\mathsf{Var}(X)+\mathbb{E}[X]^{2})(\mathsf{Var}(Y)+\mathbb{E}[Y]^{2})+\mathsf{Cov}(X^{2},Y^{2})$
@@
$\text{b. \hskip 5.69046pt}\mathsf{Var}(X+Y)=\mathsf{Var}(X)+\mathsf{Var}(Y)+2\mathsf{Cov}(X,Y),$
where the second property can be generalized for $k$ random variables $\{X_{i}\}_{i=0}^{k}$ as
@@
where the covariances vanishes based on property 1. Thus, the following equality holds
@@
This results in $\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j}s^{\iota}|_{i-j})=\mathsf{Var}(b_{\mu}(\iota)e^{\mu}|_{j})(\mathsf{Var}(s^{\iota}|_{i-j})+\mathbb{E}[s^{\iota}|_{i-j}]^{2})$. In addition, for a random variable $X$, it holds that $\mathsf{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}$. Therefore,
@@
The expression for the variance of the coefficients of the critical quantity simply follows from the observation that the covariance between $a_{\iota_{1}}s^{\iota_{1}}|_{i}$ and $a_{\iota_{2}}s^{\iota_{2}}|_{i}$ vanishes whenever $\iota_{1}\neq\iota_{2}$. Indeed, by bilinearity of the covariance operator, $\mathsf{Cov}(a_{\iota_{1}}s^{\iota_{1}}|_{i},a_{\iota_{2}}s^{\iota_{2}}|_{i})$ can be expressed as a sum of terms of the form
$\mathsf{Cov}(b_{\mu_{1}}({\iota_{1}})|_{j}(e^{\mu_{1}}s^{\iota_{1}})|_{i-j},b_{\mu_{2}}({\iota_{2}})|_{j}(e^{\mu_{2}}s^{\iota_{2}})|_{i-j})\,$,
each of which vanishes by Property 1. Therefore, we obtain
$\mathsf{Var}(\nu|_{i})=\mathsf{Var}(\sum_{\iota\geq 0}a_{\iota}s^{\iota}|_{i})=\sum_{\iota\geq 0}\mathsf{Var}(a_{\iota}s^{\iota}|_{i})+\sum_{\iota_{1}\neq\iota_{2}}\mathsf{Cov}(a_{\iota_{1}}s^{\iota_{1}}|_{i},a_{\iota_{2}}s^{\iota_{2}}|_{i})\,$,
thus concluding the proof.
Appendix 0.C On the Negligibility of Key Switching
In this work, we chose to omit and not explicitly address the contribution of key switching in our analysis of noise variance, despite its necessity for ensuring the practicality of evaluated circuits. This decision is justified by the fact that, in general, the parameters associated with key switching can be selected so that its impact on the overall noise remains negligible in comparison to that of other operations — especially the multiplication operation, which immediately precedes the key switching step.
While our approach can easily accommodate this contribution, we opted to exclude it from our presentation for the sake of clarity and due to the aforementioned reason. Various key switching methods exist. To support the reasonableness of our choice, we provide a justification of its negligible impact for one of the main variants: GHS (Gentry Halevi Smart).
Let $q$ be the modulus of the ciphertext to be relinearized, and let $Q$ be a modulus specifically chosen for this purpose, such that $q<Q$ and $q\mid Q$.
The core idea of the GHS variant is to perform relinearization in the larger ring $\mathcal{R}_{Q}$, and then apply modulus switching to return to the smaller ring $\mathcal{R}_{q}$, reducing the error.
Let $c=(c_{0},c_{1},c_{2})$ be the result of the multiplication of two ciphertexts, with $c_{0},c_{1},c_{2}\in\mathcal{R}_{q}$.
The key-switching key for the GHS variant is defined as
$({ek_{0}},{ek_{1}})=\left(\left(-a\cdot s+te+\frac{Q}{q}s^{2}\right),a\right)\mod Q\,.$
The ciphertext resulting from relinearization in $R_{Q}$ is given by
[…]
Then, this new ciphertext is scaled back modulo $q$ using the modulus switching technique previously introduced, thereby obtaining
[…]
It is easy to verify that the critical quantity of the new relinearized ciphertext is given by
$\nu_{\mathsf{ks}}=[\hat{c_{0}}+\hat{c_{1}}\cdot s]_{q}=\nu_{\mathsf{mul}}+\left[\frac{q}{Q}\left(tc_{2}\cdot e+\delta_{0}+\delta_{1}\cdot s\right)\right]_{q}\,.$
A typical choice for the size of $Q$ is $Q\approx q^{2}$ [30]. In this way, the variance of the second component will be
$V_{\frac{q}{Q}(tc_{2}\cdot e+\delta_{0}+\delta_{1}\cdot s)}=\frac{1}{q^{2}}\left({nt^{2}V_{e}V_{c_{2}}}+{V_{\delta_{0}}+nV_{s}V_{\delta_{1}}}\right)=\frac{t^{2}}{12}\left({nV_{e}}+{1+nV_{s}}\right)\,.$
On the other hand, in accordance with the circuits proposed in the paper, we may assume that the ciphertext to be relinearized results from the product of two terms that have already undergone modulus switching.
Consequently, the variance of the first component, namely $\nu_{\mathsf{mul}}$, is at least $nV_{\mathsf{ms}}^{2}F_{s}(1,1)$ where $V_{\mathsf{ms}}=\frac{t^{2}nV_{s}}{12}$, thereby making evident the negligible contribution of the key switching step.